A charge of +7 micro-coulombs is placed at x = 0 cm, y = +17 cm, a charge of -4 micro-coulombs is placed at x = 0 cm, y = -24 cm, a charge of +6 micro-coulombs is placed at x = -7 cm, y = 0 cm, and an unknown charge is placed at at x = +20 cm, y = 0 cm. The angle of the total electric field at the origin is 293 degrees counter-clockwise from the positive x axis. What is the unknown charge in micro-coulombs? If negative, include a negative sign in your answer.
Answer #1

E1 17cm o 7cm q3-6uC 20cm q4 24cm E2 q2--4uC

E=kq/r^{2}

Give θ 293°

tan293 = EyEl

----------------------

E_{y}=E_{1}+E_{2}

E_{y}=kq_{1}/r_{1}^{2}+kq_{2}/r_{2}^{2}

E_{y}=-9*10^{9}*7*10^{-6}C/(0.17m)^{2}-9*10^{9}*4*10^{-6}C/(0.24m)^{2}

E,ー-9 * 10° * 7 * 10-6/0.172-9 * 10 * 4 * 10-6/0.24°

{color{Red} E_{y}=-2.8*10^{6}N/C}

===================

tan293 = EyEl

E_{x}=E_{y}/tan293

E_{x}=(-2.8*10^{6}N/C)/tan293

{color{Red} E_{x}=1.19*10^{6}N/C}

---------------

E_{x}=E_{3}+E_{4}

1.19*10^{6}N/C=kq_{3}/r_{3}^{2}+kq_{4}/r_{4}^{2}

1.19*10^{6}N/C=9*10^{9}*6*10^{-6}C/(0.07m)^{2}-9*10^{9}*q_{4}*10^{-6}/(0.2m)^{2}

1.19*10^{6}=9*10^{3}*(6/0.07^{2}-q_{4}/0.2^{2})

132.22=1224.49-q_{4}/0.2^{2}

-1092.27=-q_{4}/0.2^{2}

ANSWER: {color{Red} q_{4}=43.69uC}

========================